Suppose your annual income is $43,194 and there are two tax brackets: 10% for $0-$20,000 and 30% for everything above $20,000.

What percentage of your income will you pay in taxes? Enter your answer in decimal (0.07 for 7%).
Answer #1

Tax on first 20000 = 20000 * 10% = 2000

tax on remaining income = (43194 - 20000) * 30%

= 6958.2

total tax paid = 8958.2

percentage of income = 8958.2/43194

= 0.2074

= 0.21